# Identifiable Deep Latent Variable Models for MNAR Data

### 25 Mar 2026 v1

@@ 2 Ground-truth model of full-data distribution
In this section, we present the ground-truth model for the distribution of all variables and missing indicators. We begin by introducing the notation. We use capital letters to denote random variables or vectors, and lowercases for their values. Consider a random vector $X=(X_{1},\cdots,X_{p})^{T}\in\mathcal{X}$, where $\mathcal{X}$ is a $d$-dimensional manifold embedded in $\mathbb{R}^{p}$ for $d\leq p$. Let $R=(R_{1},\cdots,R_{p})^{T}\in\{0,1\}^{p}$ be a vector of missingness indicators of $X$, where $1$ corresponds to “observed”, and 0 corresponds to “missing”. Let $(r^{(1)},x^{(1)}),\cdots,(r^{(n)},x^{(n)})$ denote $n$ independently and identically distributed (i.i.d.) samples of the pair $(R,X)$. With a slight abuse of notation, we write $R=r$ as a shorthand for $(R_{1},..,R_{p})=(r_{1},\cdots,r_{p})$. Let $X_{(r)}$ be a subvector of observed elements in $X$ when $R=r$. The observed data consist of i.i.d. realizations of the vector $(R,X_{(R)})$. For convenience, we define another indicator M=1RM=1-R as the opposite of $R$, that is, $M_{j}$ = 1 represents $X_{j}$ missing and 0 represents $X_{j}$ observed.
@@ 3 Conditional no self-censoring given latent variables
where $A_{-j}=(A_{1},...,A_{j-1},A_{j+1},...,A_{p})$ denotes a subvector of $A$ with the $j^{\text{th}}$ entry removed.
@@ 6.1 Imputation and generation
For higher dimensions, we generated $X$ from latent variables, experimenting with data dimensions of $p=30,50,100$ while keeping the latent dimension fixed at 10. All other parameters and settings were consistent with the previous $3$-dimensional experiments, with no latent variable used for generating missingness and with the mapping function $h$ being linear. The missing rates were controlled to be around $50\%$. In these higher-dimensional simulations, we considered different settings with block-wise missingness, where variables within the same group are either missing together or observed together. See Figure 3 for an illustration of a dataset with $3$ groups. Specifically, we varied the sizes of the groups. The term “no self-censoring” in the context of block-wise missingness refers to the scenario where the missingness of the variables in a group does not depend on any of the variables within that group. The imputation results are presented in Table 2. Due to the prohibitive computational time of the MissForest method, we excluded it from these higher-dimensional simulations. However, we retained MICE as a baseline for comparison, given its ignorability assumption in contrast to other methods.
@@
Table 2: Imputation RMSE of simulated data with higher dimensions and $10$ latent variables, where $p$ is the number of variables in $X$ and $gs$ is the group size of the variables, i.e., the number of variables within a group.
@@ 7.2 HIV positive mothers in Botswana
We learn the distribution of the data using our method with the Bernoulli distribution for the data generative model. As the first step in estimating the relationship of interest, we obtained the joint distribution using our generative model, and compared the results with other estimation methods including complete case analysis (CC), MICE, and estimation using the augmented inverse probability weighted (AIPW) by Malinsky et al. (2021) which assumes a no self-censoring model. We can see from the joint distribution in Table 4 that certain probabilities are similar for the proposed method, AIPW and MICE, which all have substantial differences from those for the complete case analysis.
@@ 7.3 Yahoo! R3 music ratings
In this subsection, we apply our method to the Yahoo! R3 dataset (Marlin and Zemel, 2009; Wang et al., 2018). This dataset is primarily designed for the analysis of user-song ratings, with a specific focus on evaluating MNAR imputation techniques. Within this dataset, we have a training set that consists of more than 300,000 self-selected ratings on a scale from $1$ to $5$ provided by 15,400 users for a selection of 1,000 songs, which has an entry-wise missing rate of around $98.0\%$. In addition, there is an MCAR test set, which comprises ratings from 5,400 users across a set of 10 songs chosen at random. We train our model on the training set and subsequently evaluate the imputation performance on the test set. We acknowledge that this dataset likely involves self-censoring. Our aim here is not to assume otherwise, but to demonstrate the robustness and practical effectiveness of our method on a complex real-world dataset.
@@
We compare the proposed method with GINA (Ma and Zhang, 2021) and not-MIWAE (Ipsen et al., 2020), which are also designed for MNAR but did not address the issue of identifiability without resorting to auxiliary variables. We model the missingness mechanism as a linear function in all methods, allowing each $R_{j}$ to depend on all dimensions of $X$ in the baselines. We did not include classical methods such as MICE or missForest due to their prohibitively long training time. We visualize the imputation results in the five categorical bar graphs, together with the training and test data. First, we can notice that the training data have a very different distribution from the test data, where the latter is known to be a relatively unbiased representation of the true data distribution, while the training data are MNAR and thus biased. More importantly, we find that all the generative models learn a data distribution closer to the test data than to the training data, where the proposed model produces the most pronounced difference between ratings $1$ and $2$ among all the methods, relatively closest to the distribution of the unbiased test data.
@@ A.6 Proof of Corollary 6
where the second term is deterministic and is either 0 or $1$, and it will always be zero for $K$ large enough according to Proposition 1. Applying Chebyshev’s inequality to the first term and setting $\tau=\frac{\epsilon}{2\sqrt{\text{Var}(\hat{\mathcal{L}}_{K})}}$, we have
@@ A.8 Proof of Theorem 8
Denote the dimension of the observed data as $p_{(r)}$ for $1\leq p_{(r)}\leq p$, with $p_{(r)}$ being the number of $1$’s in $r$. Define the function $F:\mathbb{R}^{p_{(r)}}\mapsto[0,1]^{p_{(r)}}$ as
@@ Appendix C More Details on the Choice of the Latent Dimension $\kappa_{2}$
We experimented with the choice of latent dimensions for the missingness mechanism using the UCI datasets. We simulated missingness on the original complete UCI data according to no self-censoring with the latent $q_{2}$ dimensions chosen from $0,\lfloor p/2\rfloor$ and $p$, where $p$ is the data dimension of each dataset. In each of these simulation settings, (fixing the latent dimensions for the data, which have been known for the complete datasets through cross-validation using the generative models,) we experimented with different latent dimensions for the missingness mechanism, $\kappa_{2}$ in the learning model. See Figure 7. There does not appear to be significant difference among different choices of $\kappa_{2}$ in the learning model in each case, and choosing the dimension $\kappa_{2}$ to be $1$ works sufficiently well in the learning model from these experiments, no matter what the true latent dimensions are. This empirical result is consistent with the theoretical conclusion in Lemma 7.
@@ Appendix E Implementation Details in Experiments
Prior to training, all user ratings are normalized to fall within the range of 0 to 1 (this normalization will be reversed during evaluation). Gaussian likelihood with a variance of $\gamma=0.02$ is employed for all the generative models. A 20-dimensional latent space for the data is utilized, and the decoder follows a structure of 20-10-$p$. The inference network adopts the point net structure proposed by Ma et al. (2018), employing a 20-dimensional feature mapping $h$ parameterized by a single-layer neural network and 20-dimensional ID vectors for each variable. The symmetric operator is defined as the summation operator. The missing model is parameterized by a linear neural network. All methods underwent training for 100 epochs with a batch size of 100. Imputation RMSE was estimated using 10 importance samples.

### 27 Mar 2026 v2

@@ 2 Ground-truth model of full-data distribution
In this section, we present the ground-truth model for the distribution of all variables and missing indicators. We begin by introducing the notation. We use capital letters to denote random variables or vectors, and lowercases for their values. Consider a random vector $X=(X_{1},\cdots,X_{p})^{T}\in\mathcal{X}$, where $\mathcal{X}$ is a $d$-dimensional manifold embedded in $\mathbb{R}^{p}$ for $d\leq p$. Let $R=(R_{1},\cdots,R_{p})^{T}\in\{0,1\}^{p}$ be a vector of missingness indicators of $X$, where 11 corresponds to “observed”, and 0 corresponds to “missing”. Let $(r^{(1)},x^{(1)}),\cdots,(r^{(n)},x^{(n)})$ denote $n$ independently and identically distributed (i.i.d.) samples of the pair $(R,X)$. With a slight abuse of notation, we write $R=r$ as a shorthand for $(R_{1},..,R_{p})=(r_{1},\cdots,r_{p})$. Let $X_{(r)}$ be a subvector of observed elements in $X$ when $R=r$. The observed data consist of i.i.d. realizations of the vector $(R,X_{(R)})$. For convenience, we define another indicator M=1RM=1-R as the opposite of $R$, that is, $M_{j}$ = 1 represents $X_{j}$ missing and 0 represents $X_{j}$ observed.
@@ 3 Conditional no self-censoring given latent variables
where $A_{-j}=(A_{1},...,A_{j-1},A_{j+1},...,A_{p})$ denotes a subvector of AA with the $j^{\text{th}}$ entry removed.
@@ 6.1 Imputation and generation
For higher dimensions, we generated $X$ from latent variables, experimenting with data dimensions of $p=30,50,100$ while keeping the latent dimension fixed at 10. All other parameters and settings were consistent with the previous 33-dimensional experiments, with no latent variable used for generating missingness and with the mapping function $h$ being linear. The missing rates were controlled to be around $50\%$. In these higher-dimensional simulations, we considered different settings with block-wise missingness, where variables within the same group are either missing together or observed together. See Figure 3 for an illustration of a dataset with 33 groups. Specifically, we varied the sizes of the groups. The term “no self-censoring” in the context of block-wise missingness refers to the scenario where the missingness of the variables in a group does not depend on any of the variables within that group. The imputation results are presented in Table 2. Due to the prohibitive computational time of the MissForest method, we excluded it from these higher-dimensional simulations. However, we retained MICE as a baseline for comparison, given its ignorability assumption in contrast to other methods.
@@
Table 2: Imputation RMSE of simulated data with higher dimensions and $10$ latent variables, where $p$ is the number of variables in $X$ and gsgs is the group size of the variables, i.e., the number of variables within a group.
@@ 7.2 HIV positive mothers in Botswana
We learn the distribution of the data using our method with the Bernoulli distribution for the data generative model. As the first step in estimating the relationship of interest, we obtained the joint distribution using our generative model, and compared the results with other estimation methods including complete case analysis ($C$), MICE, and estimation using the augmented inverse probability weighted (AIPW) by Malinsky et al. (2021) which assumes a no self-censoring model. We can see from the joint distribution in Table 4 that certain probabilities are similar for the proposed method, AIPW and MICE, which all have substantial differences from those for the complete case analysis.
@@ 7.3 Yahoo! R3 music ratings
In this subsection, we apply our method to the Yahoo! R3 dataset (Marlin and Zemel, 2009; Wang et al., 2018). This dataset is primarily designed for the analysis of user-song ratings, with a specific focus on evaluating MNAR imputation techniques. Within this dataset, we have a training set that consists of more than 300,000 self-selected ratings on a scale from 11 to 55 provided by 15,400 users for a selection of 1,000 songs, which has an entry-wise missing rate of around $98.0\%$. In addition, there is an MCAR test set, which comprises ratings from 5,400 users across a set of 10 songs chosen at random. We train our model on the training set and subsequently evaluate the imputation performance on the test set. We acknowledge that this dataset likely involves self-censoring. Our aim here is not to assume otherwise, but to demonstrate the robustness and practical effectiveness of our method on a complex real-world dataset.
@@
We compare the proposed method with GINA (Ma and Zhang, 2021) and not-MIWAE (Ipsen et al., 2020), which are also designed for MNAR but did not address the issue of identifiability without resorting to auxiliary variables. We model the missingness mechanism as a linear function in all methods, allowing each $R_{j}$ to depend on all dimensions of $X$ in the baselines. We did not include classical methods such as MICE or missForest due to their prohibitively long training time. We visualize the imputation results in the five categorical bar graphs, together with the training and test data. First, we can notice that the training data have a very different distribution from the test data, where the latter is known to be a relatively unbiased representation of the true data distribution, while the training data are MNAR and thus biased. More importantly, we find that all the generative models learn a data distribution closer to the test data than to the training data, where the proposed model produces the most pronounced difference between ratings 11 and 22 among all the methods, relatively closest to the distribution of the unbiased test data.
@@ A.6 Proof of Corollary 6
where the second term is deterministic and is either 0 or 11, and it will always be zero for $K$ large enough according to Proposition 1. Applying Chebyshev’s inequality to the first term and setting $\tau=\frac{\epsilon}{2\sqrt{\text{Var}(\hat{\mathcal{L}}_{K})}}$, we have
@@ A.8 Proof of Theorem 8
Denote the dimension of the observed data as $p_{(r)}$ for $1\leq p_{(r)}\leq p$, with $p_{(r)}$ being the number of 11’s in $r$. Define the function $F:\mathbb{R}^{p_{(r)}}\mapsto[0,1]^{p_{(r)}}$ as
@@ Appendix C More Details on the Choice of the Latent Dimension $\kappa_{2}$
We experimented with the choice of latent dimensions for the missingness mechanism using the UCI datasets. We simulated missingness on the original complete UCI data according to no self-censoring with the latent $q_{2}$ dimensions chosen from $0,\lfloor p/2\rfloor$ and $p$, where $p$ is the data dimension of each dataset. In each of these simulation settings, (fixing the latent dimensions for the data, which have been known for the complete datasets through cross-validation using the generative models,) we experimented with different latent dimensions for the missingness mechanism, $\kappa_{2}$ in the learning model. See Figure 7. There does not appear to be significant difference among different choices of $\kappa_{2}$ in the learning model in each case, and choosing the dimension $\kappa_{2}$ to be 11 works sufficiently well in the learning model from these experiments, no matter what the true latent dimensions are. This empirical result is consistent with the theoretical conclusion in Lemma 7.
@@ Appendix E Implementation Details in Experiments
Prior to training, all user ratings are normalized to fall within the range of 0 to 1 (this normalization will be reversed during evaluation). Gaussian likelihood with a variance of $\gamma=0.02$ is employed for all the generative models. A 20-dimensional latent space for the data is utilized, and the decoder follows a structure of 20-10-pp. The inference network adopts the point net structure proposed by Ma et al. (2018), employing a 20-dimensional feature mapping $h$ parameterized by a single-layer neural network and 20-dimensional ID vectors for each variable. The symmetric operator is defined as the summation operator. The missing model is parameterized by a linear neural network. All methods underwent training for 100 epochs with a batch size of 100. Imputation RMSE was estimated using 10 importance samples.